OpenStudy (anonymous):

sand falling onto a conical pile at a rate of 10 cubic feet/min. diameter of the base of the cone is 3times the altitude. at what rate is the height of the pile changing when the pile is 15ft high?

OpenStudy (anonymous):

we know V= pi/3 r^2 *h also given 2r =3h hence r =3h/2 (assume r and h to be fucnction of time t and hence V is also function of t) thus we have V= pi/3 *9*h^2/4 *h or V=3pi/4 *h^3 now diiferentiating both side wrt t we have dV/dt =9 pi/4 * h^2 dh/dt here given dV/dt = 10 cubic feet/min and h =15 we need to find dh/dt